### v1

# Approximation schemes for covering and packing mixed-integer programs with a fixed number of constraints

Abstract
This paper presents an algorithmic study of a class of covering mixed-integer linear programming problems which encompasses classic cover problems, including multidimensional knapsack, facility location and supplier selection problems. We first show some properties of the vertices of the associated polytope, which are then used to decompose the problem into instances of the multidimensional knapsack cover problem with a single continuous variable per dimension. The proposed decomposition is used to design a polynomial-time approximation scheme for the problem with a fixed number of constraints. To the best of our knowledge, this is the first approximation scheme for such a general class of covering mixed-integer programs. Moreover, we design a fully polynomial-time approximation scheme and an approximate linear programming formulation for the case with a single constraint. These results improve upon the previously best-known 2-approximation algorithm for the knapsack cover problem with a single continuous variable. Finally, we show a perfect compact formulation for the case where all variables have the same lower and upper bounds. Analogous results are derived for the packing and assignment variants of the problem.
1 Introduction
This paper presents an algorithmic study of mixed-integer programming problems, denoted by p, which are defined as:
$\min\;\;$ $\sum_{i\in[n]}\sum_{j\in[m]}v_{ij}x_{ij}+\sum_{i\in[n]}f_{i}y_{i}$ (1)
s.t. $\sum_{i\in[n]}x_{ij}\geq d_{j}$ $\forall j\in[m],$ (2)
$\ell_{ij}y_{i}\leq x_{ij}\leq c_{ij}y_{i}$ $\forall i\in[n],\;j\in[m],$ (3)
$y_{i}\in\{0,1\}$ $\forall i\in[n].$ (4)
where $m$ and $n$ are positive integers, $v,\ell,c\in\mathbb{Z}_{\geq}^{nm}$, $d\in\mathbb{Z}_{\geq}^{m}$ and $f\in\mathbb{Z}_{\geq}^{n}$ denote the nonnegative integer parameters and $a]=\{1,\ldots,a\}$ for a positive integer $a$. Throughout this paper, it is assumed, without loss of generality, that $d_{j}\geq c_{ij}\geq\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$. For completeness, we also consider the packing variant of p, defined analogously as:
$\max\;\;$ $\sum_{i\in[n]}\sum_{j\in[m]}v_{ij}x_{ij}+\sum_{i\in[n]}f_{i}y_{i}$
s.t. $\sum_{i\in[n]}x_{ij}\leq d_{j}$ $\forall j\in[m],$
$\ell_{ij}y_{i}\leq x_{ij}\leq c_{ij}y_{i}$ $\forall i\in[n],\;j\in[m],$
$y_{i}\in\{0,1\}$ $\forall i\in[n].$
under the same assumptions on all parameters.
The covering integer programs (cip) are closely related to p. Given a matrix $A\in\mathbb{R}^{m\times n}_{\geq}$, vectors $a\in\mathbb{R}^{m}_{\geq}$, and $b,h\in\mathbb{R}^{n}_{\geq}$, cip is defined as $\min\{h^{T}z\mathrel{\mathop{\ordinarycolon}}Az\geq a,z\leq b,z\in\mathbb{Z}^{n}_{\geq}\}$. carr2000strengthening propose a $\Delta_{1}$-approximation for this problem, where $\Delta_{1}$ is the maximum number of non-zero coefficients in any row of $A$. kolliopoulos2005approximation design a $\mathcal{O}(\log m)$-approximation for cip, and observe that this is asymptotically the best possible unless $\mathsf{P}=\mathsf{NP}$. This inapproximability threshold follows from the classic set cover problem, which does not admit a $o(\log m)$-approximation unless $\mathsf{P}=\mathsf{NP}$ (raz1997sub). The set cover problem is a special case of cip where $z$ is binary, and binary cip is a particular case of p where $\ell_{ij}=c_{ij}=A_{ij}$, $v_{ij}=0$, and $f_{i}=h_{i}$ for all $i\in[n]$ and $j\in[m]$. As a consequence, p does not admit a $o(\log m)$-approximation unless $\mathsf{P}=\mathsf{NP}$. Although there is little hope of designing better approximations for p in general, the case where the number of constraints is fixed remains largely unexplored. Observe that p with fixed $m$ encompasses classic cover problems, including multidimensional knapsack, facility location and supplier selection problems. Therefore, p remains $\mathsf{NP}$-hard even when $m=1$. For any fixed $m\geq 2$p with $m$ constraints does not admit a fully polynomial-time approximation scheme (FPTAS) unless $\mathsf{P}=\mathsf{NP}$. This is due to the analogous hardness result for multidimensional knapsack cover that can be obtained following essentially the same proof for the multidimensional knapsack (packing) devised by magazine1984note, and to the fact that multidimensional knapsack cover is a particular case of p with a fixed number of constraints. On the positive side, friezeApproximationAlgorithmsMdimensional1984 design a polynomial-time approximation scheme (PTAS) to the multidimensional knapsack cover problem. For the knapsack cover problem (i.e., the one-dimensional case), guntzer2000approximate propose an FPTAS which is based on a greedy strategy.
The packing integer programs (pip) are defined analogously and encompass classic combinatorial optimization problems such as knapsack, matching, and independent set problems. As a consequence, binary pip (and thus the packing variant of p) cannot be approximated within $n^{1-\epsilon}$, for any $\epsilon>0$, unless $\mathsf{P}=\mathsf{NP}$ (zuckerman2006linear). There exists a vast literature on approximation algorithms for binary pip, nearly all of which rely on randomized rounding techniques (see, e.g., srinivasan1995improved; bansal2012solving; chekuri2004multidimensional).
[…]
For p with a single constraint, we devise an FPTAS based on a dynamic programming approach in Section 4. A byproduct of this result is an FPTAS for the knapsack cover with a single continuous variable, thereby improving upon the previously best-known 2-approximation algorithm by zhaoApproximationAlgorithms012014. We note that vanhoeselFullyPolynomialApproximation2001 present an FPTAS for the single-item capacitated economic lot-sizing problem that is more general than one-dimensional p in its consideration of backlogging and holding costs. However, their model does not consider lower bounds. Additionally, for each $\epsilon\in(0,1)$, we prove the existence of a compact linear programming formulation which yields a solution of value at most $(1+\epsilon)$ times the optimal value of the problem.
Finally, in Section 5, we show a perfect compact formulation for the one-dimensional p with uniform bounds, that is, $\ell_{i}=\ell$ and $c_{i}=c$ for all $i\in[n]$. constantinoLowerBoundsLotSizing1998a investigates the uncapacitated version of this problem, i.e., the case where $c\geq d$. More precisely, he derives two families of valid inequalities and shows that they describe the convex hull in this special case. Although one-dimensional uniform p is a special case (with positive demand only in the last period) of a single-item lot-sizing problem with piecewise linear costs, which is known to be polynomial-time solvable (hellionPolynomialTimeAlgorithm2012a), no tight formulation is known for this problem. Hence the proposed perfect compact formulation contributes in this direction. Concluding remarks and directions for further investigation are presented in Section 6. Note that each section also includes an analysis of the corresponding packing problem.
2 Decomposition into multidimensional knapsack problems
This section begins by showing a simple characterization of the instances of p with optimal value equal to zero.
Proposition 2.1.
Let $I$ be an instance of p, and let $S=\{i\in[n]\mathrel{\mathop{\ordinarycolon}}f_{i}=0\}$. It holds that the optimal value $\operatorname{OPT}(I)=0$ if and only if, for each $j\in[m]$, $\sum_{i\in Z_{j}}c_{ij}\geq d_{j}$, where $Z_{j}=\{i\in S\mathrel{\mathop{\ordinarycolon}}v_{ij}=0\}$.
Proof.
Suppose first that $\operatorname{OPT}(I)=0$, and let $(x,y)$ be an optimal solution of $I$. Clearly, $f_{i}=0$ for each $i\in[n]$ with $y_{i}=1$. Consider $i\in[n]$ with $y_{i}=1$, and $j\in[m]$. If $v_{ij}>0$, then $x_{ij}=0$. Hence, $\sum_{i\in Z_{j}}c_{ij}\geq\sum_{i\in Z_{j}}x_{ij}=\sum_{i\in[n]}x_{ij}\geq d_{j}$.
For the converse, define a vector $(x,y)\in\mathbb{R}^{nm}\times\{0,1\}^{n}$ that equals zero except for the entries $y_{i}=1$ for all $i\in[n]$ with $f_{i}=0$, and $x_{ij}=c_{ij}$ for all $j\in[m]$ and $i\in Z_{j}$. If follows from the construction that $(x,y)$ is a feasible solution of $I$ of value equal to zero. Hence, $\operatorname{OPT}(I)=0$. ∎
Using the previous result, one may recognize and solve any instance with optimal value equal to zero in $\mathcal{O}(nm)$ time. Henceforth, we assume that all considered instances of p have a positive optimal value.
We next present properties of an optimal solution of p. These allow us to demonstrate how to decompose p into a collection of instances of the following generalization of the multidimensional knapsack cover problem.
Multidimensional knapsack cover problem with one continuous variable per dimension (mkc).
Input: $\eta,\mu\in\mathbb{Z}_{>}$, $\bar{f}\in\mathbb{Z}^{\eta}_{\geq}$, $\bar{v}\in\mathbb{Z}^{\mu}_{\geq}$, $\bar{c}\in\mathbb{Z}^{\mu}_{>}$, $\bar{w}\in\mathbb{Z}^{\eta\mu}_{\geq}$, and $\bar{d}\in\mathbb{Z}^{\mu}_{\geq}$.
Output: Subset $S\subseteq[\eta]$, and nonnegative $\alpha_{j}\leq\bar{c}_{j}$ for all $j\in[\mu]$ such that $\sum_{i\in S}\bar{w}_{ij}\geq d_{j}-\alpha_{j}$ for every $j\in[\mu]$.
Objective: Minimize $\sum_{i\in S}\bar{f}_{i}+\sum_{j\in[\mu]}\bar{v}_{j}\alpha_{j}$.
Note that an instance of mkc is feasible if and only if $\sum_{i\in[\eta]}\bar{w}_{ij}\geq d_{j}-\bar{c}_{j}$ for every $j\in[\mu]$. We assume henceforth that all considered instances of mkc are feasible.
[…]
Lemma 2.2.
There exists an optimal solution $(x,y)$ to p such that the following assertions hold for every $j\in[m]$.
  1. (i)
    There exists at most one $g(j)\in[n]$ such that $x_{g(j),j}\in(\ell_{g(j),j},c_{g(j),j})$. Precisely, $\ell_{g(j),j}<x_{g(j),j}<c_{g(j),j}$ if any exists. Otherwise, let $g(j)\in\arg\max_{i\in[n]}\{v_{ij}\mathrel{\mathop{\ordinarycolon}}x_{ij}=c_{ij},y_{i}=1\}$ if $x_{ij}\in\{\ell_{ij}y_{i},c_{ij}y_{i}\}$ for all $i\in[n]$ and there is at least one $i\in[n]$ with $x_{ij}=c_{ij}$, and $g(j)\in\arg\min_{i\in[n]}\{v_{ij}\mathrel{\mathop{\ordinarycolon}}x_{ij}=\ell_{ij},y_{i}=1\}$ otherwise. We define $g(j)$ as the fractional variable for each $j\in[m]$.
  2. (ii)
    For each $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}\neq v_{g(j),j}$, we have $x_{ij}=\ell_{ij}$ if $v_{ij}>v_{g(j),j}$, and $x_{ij}=c_{ij}$ otherwise.
  3. (iii)
    For each $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}=v_{g(j),j}$, we have $x_{ij}=\ell_{ij}$ if $i<g(j)$, and $x_{ij}=c_{ij}$ otherwise.
Proof.
Let $(x,y)\in\mathbb{R}^{nm+n}$ be an optimal solution to p that is an extreme point of $conv(X)$, and let $j\in[m]$. Suppose to the contrary that $\ell_{aj}<x_{aj}<c_{aj}$ and $\ell_{bj}<x_{bj}<c_{bj}$ for some $a,b\in[n]$ with $a\neq b$. Let $\epsilon=\min_{d\in\{a,b\}}\min\{c_{dj}-x_{dj},\>x_{dj}-\ell_{dj}\}$. We define two vectors $x^{1},x^{2}\in\mathbb{R}^{nm+n}$ such that $x^{1}$ and $x^{2}$ are equal to $x$ except for the following entries: $x_{aj}^{1}=x_{aj}+\epsilon$, $x_{bj}^{1}=x_{bj}-\epsilon$, $x_{aj}^{2}=x_{aj}-\epsilon$, and $x_{bj}^{2}=x_{bj}+\epsilon$. It is clear that $(x^{1},y)$ and $(x^{2},y)$ belong to $X$. Moreover, $(x,y)=\frac{1}{2}(x^{1},y)+\frac{1}{2}(x^{2},y)$, a contradiction to the choice of $(x,y)$. Hence, there exists at most one $a\in[n]$ with $\ell_{aj}<x_{aj}<c_{aj}$.
Suppose first that such $a$ exists and let $g(j)=a$, i.e., $\ell_{g(j),j}<x_{g(j),j}<c_{g(j),j}$. Suppose to the contrary that $y_{i}=1$, $x_{ij}=c_{ij}>\ell_{ij}$ and $v_{ij}>v_{g(j),j}$ for some $i\in[n]\setminus\{g(j)\}$. Define a vector $\bar{x}\in\mathbb{R}^{nm}$ to be equal to $x$ except for $\bar{x}_{ij}=x_{ij}-\epsilon$ and $\bar{x}_{g(j),j}=x_{g(j),j}+\epsilon$, where $\epsilon\mathrel{\mathop{\ordinarycolon}}=\min\{c_{g(j),j}-x_{g(j),j},\>x_{ij}-\ell_{ij}\}$. It follows that $(\bar{x},y)$ is a feasible solution to p of cost strictly smaller than $(x,y)$, a contradiction. Analogously, one can prove that $x_{ij}=\ell_{ij}$ if $y_{i}=1$ and $v_{ij}>v_{g(j),j}$. Suppose now that $x_{ij}\in\{\ell_{ij}y_{i},c_{ij}y_{i}\}$ for all $i\in[n]$. Note that $v_{ij}\geq v_{i^{\prime}j}$ for all $i,i^{\prime}\in[n]$ such that $y_{i}=y_{i^{\prime}}=1$, $x_{ij}=\ell_{ij}<c_{ij}$ and $x_{i^{\prime}j}=c_{i^{\prime}j}>\ell_{i^{\prime}j}$. Otherwise, one could obtain a solution cheaper than $(x,y)$ by increasing $x_{ij}$ and decreasing $x_{i^{\prime}j}$ by the same (sufficiently small) amount. Let $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}\neq v_{g(j),j}$. If $x_{g(j),j}=c_{g(j),j}$, then the previous observation together with the definition of $g(j)$ in (i) imply that $x_{ij}=c_{ij}$ if $v_{ij}<v_{g(j),j}$, and $x_{ij}=\ell_{ij}$ if $v_{ij}>v_{g(j),j}$. If $x_{g(j),j}=\ell_{g(j),j}$, then there is no $i\in[n]$ with $y_{i}=1$ and $x_{ij}=c_{ij}$. Thus (ii) follows from the choice of $g(j)$.
Consider now $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}=v_{g(j),j}$. We define the set of inversions of any solution $(x^{\prime},y^{\prime})$ satisfying (i) as
$\{i\in\{1,\ldots,g(j)-1\}\mathrel{\mathop{\ordinarycolon}}x^{\prime}_{ij}=c_{ij},y^{\prime}_{i}=1\}\cup\{i\in\{g(j)+1,\ldots,n\}\mathrel{\mathop{\ordinarycolon}}x^{\prime}_{ij}=\ell_{ij},y^{\prime}_{i}=1\}.$
Let $V$ be the set of inversion of $(x,y)$. Suppose that $(x,y)$ has at least one inversion (i.e., $V\neq\emptyset$). Assume, without loss of generality, that this inversion is due to an element in $[n]$ smaller than $g(j)$, and choose $i\in V$ to be the largest number such that $i<g(j)$. For each $i^{\prime}\in\{i+1,\ldots,g(j)\}$ with $y_{i^{\prime}}=1$ and $v_{i^{\prime}j}=v_{g(j),j}$, starting from $g(j)$ to $i+1$, we first move from $x_{ij}$ to $x_{i^{\prime}j}$ the maximum quantity $\epsilon$ such that either $x_{ij}-\epsilon=\ell_{ij}$ or $x_{i^{\prime}j}+\epsilon=c_{i^{\prime}j}$. Then we are done in the former case, or we repeat the procedure with $i^{\prime}-1$ or end the procedure when $i^{\prime}-1=i$ in the latter one. Let $\bar{x}\in\mathbb{R}^{nm}$ be the vector obtained at the end of this procedure. Note that there is $t\in\{i,i+1,\ldots,g(j)\}$ with $y_{t}=1$ and $v_{tj}=v_{g(j),j}$ such that, for each $i^{\prime}\in\{i,\ldots,g(j)\}$ with $y_{i^{\prime}}=1$ and $v_{i^{\prime}j}=v_{g(j),j}$, we have $\bar{x}_{i^{\prime}j}=\ell_{i^{\prime}j}$ if $i^{\prime}<t$ and $\bar{x}_{i^{\prime}j}=c_{i^{\prime}j}$ if $i^{\prime}>t$. Hence, the number of inversions of $(\bar{x},y)$ is strictly smaller than the number of inversions of $(x,y)$. Moreover, since $v_{i^{\prime}j}=v_{g(j),j}$ for all $i^{\prime}\in V$, it is clear that $(\bar{x},y)$ is an optimal solution to p that still respects (i) and (ii). By repeating the previous procedure at most n1n-1 times, we obtain an optimal solution to p which satisfies (i), (ii), and (iii). The case of an inversion involving an element in $[n]$ larger than $g(j)$ is handled analogously.
Let $j\in[m]$ and $k\in[n]$. We define $L_{kj}=\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}>v_{kj}\}\cup\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}=v_{kj}\text{ and }i<k\}$ and $C_{kj}=\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}<v_{kj}\}\cup\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}=v_{kj}\text{ and }i>k\}.$ Note that $\{L_{kj},C_{kj}\}$ is a partition of $n]\setminus\{k\}$ which only depends on the input instance, and can be easily computed in linear time.
Theorem 2.3.
Let $I$ be an instance of p. There exists a function $g\colon[m]\to[n]$ such that $I$ admits an optimal solution $(x,y)\in\mathbb{R}^{nm}\times\{0,1\}^{n}$ where, for each $j\in[m]$, $y_{g(j)}=1$, $x_{ij}=\ell_{ij}y_{i}$ for all $i\in L_{g(j),j}$ and $x_{ij}=c_{ij}y_{i}$ for all $i\in C_{g(j),j}$.
Proof.
The proof is immediate from Lemma 2.2 and the definitions of the $L$ and $C$ sets. ∎
It follows from Theorem 2.3 that problem p boils down to finding such a function $g$ and solving the mkc instance associated with $g$. Using this idea, we next propose an alternative formulation for p applied to the (sub)partitions of $[n]$ given by $L$ and $C$ for all possible $\mathcal{O}(n^{m})$ choices of $(g(1),\ldots,g(m))\in[n]^{m}$.
Let $\mathcal{G}$ be the set of all functions $g\colon[m]\to[n]$. For every $g\in\mathcal{G}$, $i\in[n]$ and $j\in[m]$, define
$w_{ij}^{g}=\begin{cases}\ell_{ij}&\text{ if $i\in L_{g(j),j}$,}\\ c_{ij}&\text{ if $i\in C_{g(j),j}$,}\\ \ell_{g(j)j}&\text{ if $i=g$.}\end{cases}$
Let $g\in\mathcal{G}$. For each $j\in[m]$, define $c^{g}_{g(j),j}=c_{g(j),j}-\ell_{g(j),j}$, and $v^{g}_{j}=v_{g(j),j}$. For each $i\in[n]$, define $f^{g}_{i}=f_{i}+\sum_{j\in[m]}v_{ij}w^{g}_{ij}$.
The following mixed-integer linear formulation has binary variables $z^{g}$ and $y^{g}_{i}$ for all $g\in\mathcal{G}$ and $i\in[n]$, and real variables $\alpha^{g}_{j}$ for all $g\in\mathcal{G}$ and $j\in[m]$.
$\min\;\;$ $\sum_{g\in G}\left(\sum_{i\in[n]}f^{g}_{i}y_{i}^{g}+\sum_{j\in[m]}v_{j}^{g}\alpha_{j}^{g}\right)$ (5)
s.t. $\sum_{g\in G}z^{g}=1,$ (6)
$y^{g}_{i}\leq z^{g}$ $\forall g\in\mathcal{G},i\in[n],$ (7)
$y^{g}_{g(j)}=z^{g}$ $\forall g\in\mathcal{G},j\in[m],$ (8)
$\sum_{i\in[n]}w^{g}_{ij}y^{g}_{i}+\alpha^{g}_{j}\geq d_{j}z^{g}$ $\forall g\in\mathcal{G},j\in[m],$ (9)
$c^{g}_{g(j),j}z^{g}\geq\alpha^{g}_{j}\geq 0$ $\forall g\in\mathcal{G},j\in[m],$ (10)
$z^{g}\in\{0,1\}$ $\forall g\in\mathcal{G},$ (11)
$y^{g}_{i}\in\{0,1\}$ $\forall g\in\mathcal{G},i\in[n].$ (12)
Proposition 2.4.
The formulation (5)–(12) correctly models p.
Proof.
Let $I$ be an instance of p. Let $g\colon[m]\to[n]$ and $(x,y)$ as given in Theorem 2.3. Let us define a vector $(z,\bar{y},\alpha)\in\{0,1\}^{|\mathcal{G}|}\times\{0,1\}^{|\mathcal{G}|n}\times\mathbb{R}^{|\mathcal{G}|m}$ with all entries equal to zero except for $z^{g}=1$, $\bar{y}^{g}=y$, and $\alpha^{g}_{j}=x_{g(j),j}-\ell_{g(j),j}$ for all $j\in[m]$. One may easily verify that $(z,\bar{y},\alpha)$ satisfies constraints (6)–(12) as $(x,y)$ satisfies (2)–(4). Thus, the objective function (5) on $(z,\bar{y},\alpha)$ is equal to
$\sum_{i\in[n]}f^{g}_{i}\bar{y}_{i}^{g}+\sum_{j\in[m]}v_{j}^{g}\alpha_{j}^{g}$ $=\sum_{i\in[n]}\sum_{j\in[m]}v_{ij}w^{g}_{ij}\bar{y}^{g}_{i}+\sum_{i\in[n]}f_{i}\bar{y}^{g}_{i}+\sum_{j\in[m]}v^{g}_{j}\alpha^{g}_{j}$
$\begin{split}&=\sum_{j\in[m]}\left(\sum_{i\in L_{g(j),j}}v_{ij}\ell_{ij}\bar{y}_{i}^{g}+\sum_{i\in C_{g(j),j}}v_{ij}c_{ij}\bar{y}_{i}^{g}+v_{g(j)j}\ell_{g(j)j}\bar{y}_{g(j)}^{g}\right)\\ &\phantom{=}+\sum_{i\in[n]}f_{i}\bar{y}^{g}_{i}+\sum_{j\in[m]}v^{g}_{j}\alpha^{g}_{j}\end{split}$ (13)
$=\sum_{j\in[m]}\sum_{i\in[n]\setminus\{g(j)\}}v_{ij}x_{ij}+\sum_{i\in[n]}f_{i}y_{i}+\sum_{j\in[m]}v_{g(j),j}x_{g(j),j}$ (14)
$=\sum_{j\in[m]}\sum_{i\in[n]}v_{ij}x_{ij}+\sum_{i\in[n]}f_{i}y_{i},$
where (LABEL:eq:w-definition) follows from the definition of $w$, and (14) holds because, for each $j\in[m]$ and $i\in[n]\setminus\{g(j)\}$, we have $x_{ij}=\ell_{ij}y_{i}$ if $i\in L_{g(j),j}$, $x_{ij}=c_{ij}y_{i}$ if $i\in C_{g(j),j}$, and $x_{ij}=\alpha^{g}_{j}+\ell_{g(j),j}$ otherwise. Finally, it is clear that any feasible solution of (6)–(12) induces a feasible solution of (2)–(4) with the same value. Therefore, (5)–(12) is a correct formulation of p. ∎
This formulation reveals a decomposition of p into a collection of mkc instances. Precisely, for each $g\in\mathcal{G}$, the formulation for $g$ is
$\min\;\;$ $\sum_{i\in[n]}f^{g}_{i}y^{g}_{i}+\sum_{j\in[m]}v^{g}_{j}\alpha^{g}_{j}$ (15)
s.t. $\sum_{i\in[n]}w^{g}_{ij}y^{g}_{i}+\alpha^{g}_{j}\geq d_{j}$ $\forall j\in[m],$ (16)
$c^{g}_{g(j),j}\geq\alpha^{g}_{j}\geq 0$ $\forall j\in[m],$ (17)
$y^{g}_{g(j)}=1$ $\forall j\in[m],$ (18)
$y^{g}_{i}\in\{0,1\}$ $\forall i\in[n].$ (19)
Note that, for every $i\in[n]$, $f^{g}_{i}$ is the cost of choosing item $i$, and $w^{g}_{ij}$ is the weight of item $i$ in knapsack $j$, where $j\in[m]$. For each $j\in[m]$, $d_{j}$ is the demand of knapsack $j$, $v^{g}_{j}$ and $c^{g}_{g(j),j}$ are the cost and the upper bound of the continuous variable associated with knapsack $j$, respectively.
We conclude this section with the following corollary of Proposition 2.4, which is used in the algorithms designed in the next sections.
Corollary 2.5.
Let $I$ be an instance of p. There exists a collection $\mathcal{I}$ of $\mathcal{O}(n^{m})$ instances of mkc such that
$\operatorname{OPT}(I)=\min_{I^{\prime}\in\mathcal{I}}\operatorname{OPT}_{\textsc{mkc}}(I^{\prime}).$
The formulation developed for p can be naturally adapted to its packing counterpart. By redefining the sets $L_{kj}=\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}<v_{kj}\}\cup\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}=v_{kj}\text{ and }i>k\}$ and $C_{kj}=\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}>v_{kj}\}\cup\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}=v_{kj}\text{ and }i<k\}$, we can decompose the problem into a collection of instances of multidimensional knapsack packing problem with one continuous variable per dimension, which is denoted by mkp.
3 A polynomial-time approximation scheme
Next, we design a polynomial-time approximation scheme for p (with $m$ fixed), which is based on the decomposition presented in the previous section. friezeApproximationAlgorithmsMdimensional1984 propose a PTAS for the Multidimensional Knapsack Cover problem without continuous variables. In what follows, we extend their PTAS to include exactly one continuous variable per dimension. The analysis of the proposed algorithm follows the methodology of friezeApproximationAlgorithmsMdimensional1984.
Let $\eta,\mu\in\mathbb{Z}_{>}$, $\bar{f}\in\mathbb{Z}^{\eta}_{\geq}$, $\bar{v}\in\mathbb{Z}^{\mu}_{\geq}$, $\bar{c}\in\mathbb{Z}^{\mu}_{>}$, $\bar{w}\in\mathbb{Z}^{\eta\mu}_{\geq}$, and $\bar{d}\in\mathbb{Z}^{\mu}_{\geq}$. The Multidimensional Knapsack Cover problem with one continuous variable per dimension (mkc) is equivalent to
$\min\;\;$ $\sum_{i\in[\eta]}\bar{f}_{i}y_{i}+\sum_{j\in[\mu]}\bar{v}_{j}\alpha_{j}$ (20)
s.t. $\sum_{i\in[\eta]}\bar{w}_{ij}y_{i}\geq\bar{d}_{j}-\alpha_{j}$ $\forall j\in[\mu],$ (21)
$\bar{c}_{j}\geq\alpha_{j}\geq 0$ $\forall j\in[\mu],$ (22)
$y_{i}\in\{0,1\}$ $\forall i\in[\eta].$ (23)
Let $\epsilon>0$, and let $k=\min\{\eta,\>\lceil\mu/\epsilon\rceil\}$. We next design an algorithm $\mathcal{A}_{\epsilon}$ for mkc that runs in polynomial time when $\mu$ (a.k.a. dimension) is fixed. For every $S\subseteq[\eta]$, define $T(S)=\{i\in[\eta]\setminus S\mathrel{\mathop{\ordinarycolon}}\bar{f}_{i}>\min_{t\in S}\bar{f}_{t}\}$. Let $LP(S)$ be the linear relaxation of the formulation obtained from (20)–(23) when replacing (23) by, for all $i\in[\eta]$, $0\leq y_{i}\leq 1$ and
$y_{i}=\begin{cases}1&\text{ if }i\in S,\\ 0&\text{ if }i\in T(S).\end{cases}$
The proposed algorithm solves $LP(S)$ for every $S\subseteq[\eta]$ such that $|S|\leq k$, and then rounds up the corresponding optimal fractional solution. Finally, it outputs the best solution found. This procedure is formally described in Algorithm 1.
Algorithm 1 Algorithm $\mathcal{A}_{\epsilon}$ for mkc.
1:An instance $I=(\eta,\mu,\bar{f},\bar{v},\bar{c},\bar{w},\bar{d})$ of mkc.
2:A feasible solution $(y,\alpha)\in\{0,1\}^{\eta}\times\mathbb{R}^{\mu}$ to $I$.
3:
4:procedure $\mathcal{A}_{\epsilon}$(II)
5:  $k$ $\leftarrow$ $\min\{\eta,\>\lceil\mu/\epsilon\rceil\}$
6:  $\zeta$ $\leftarrow$ $\infty$
7:  for $S\subseteq[\eta]$ with $|S|\leq k$ do
8:   Let $\bar{d}_{j}(S)\mathrel{\mathop{\ordinarycolon}}=\bar{d}_{j}-\bar{c}_{j}-\sum_{i\in S}w_{ij}$ for each $j\in[\mu]$
9:   if $\sum_{i\in[\eta]\setminus(S\cup T(S))}\bar{w}_{ij}\geq\bar{d}_{j}(S)$ for each $j\in[\mu]$ then $\triangleright$ Otherwise, $LP(S)$ is infeasible
10:     Compute an optimal basic solution $(y^{\prime}(S),\alpha^{\prime}(S))$ to $LP(S)$
11:     $y^{\prime\prime}_{i}(S)$ $\leftarrow$ $\left\lceil y^{\prime}_{i}(S)\right\rceil$ for every $i\in[\eta]$ $\triangleright$ Rounding up to an integer solution
12:     $\zeta(S)$ $\leftarrow$ $\sum_{i\in[\eta]}f_{i}y^{\prime\prime}_{i}(S)+\sum_{j\in[\mu]}\bar{v}_{j}\alpha^{\prime}_{j}(S)$
13:     if $\zeta>\zeta(S)$ then
14:       $\zeta$ $\leftarrow$ $\zeta(S)$,   $y$ $\leftarrow$ $y^{\prime\prime}(S)$,   $\alpha$ $\leftarrow$ $\alpha^{\prime}(S)$
15:  return $(y,\alpha)$
Theorem 3.1.
Let $\epsilon>0$. Algorithm 1 is a $(1+\epsilon)$-approximation for mkc with fixed dimension.
Proof.
The proof is very similar to the proof in friezeApproximationAlgorithmsMdimensional1984. We give it here for completeness.
Let $I=(\eta,\mu,\bar{f},\bar{v},\bar{c},\bar{w},\bar{d})$ be an instance of mkc, and let $(y,\alpha)$ be the output of Algorithm 1 on $I$ (i.e., $(y,\alpha)=\mathcal{A}_{\epsilon}(I)$). Line 9 guarantees that $LP(S)$ has an optimal solution, say $(y^{\prime}(S),\alpha^{\prime}(S))$, where $S\subseteq[\eta]$ with $|S|\leq k$. By rounding up the entries of $y^{\prime}(S)$ in line 11, we obtain a vector satisfying constraints (21)–(23). Thus, the algorithm produces a solution to $I$. Let us denote by $\zeta$ the value of $(y,\alpha)$. Consider an optimal solution $(y^{*},\alpha^{*})$ to $I$, and denote by $\zeta^{*}$ the optimal value of $I$. Let us define $S^{*}=\{i\in[\eta]\mathrel{\mathop{\ordinarycolon}}y_{i}^{*}=1\}$. If $|S^{*}|\leq k$, then we have $\zeta\leq\zeta(S^{*})\leq\zeta^{*}$ as $\bar{f}$ is nonnegative, which implies $\zeta=\zeta^{*}$. Since the condition $|S^{*}|\leq k$ is always satisfied when $k=\eta$, we henceforth assume $k=\lceil\mu/\epsilon\rceil$. If $|S^{*}|>k$, then consider an ordering of $S^{*}=\{i(1),\ldots,i(r)\}$ such that $r\mathrel{\mathop{\ordinarycolon}}=|S^{*}|$ and $\bar{f}_{i(1)}\geq\ldots\geq\bar{f}_{i(r)}$. We define $S_{k}^{*}=\{i(1),\ldots,i(k)\}$ and $\sigma=\sum_{t\in[k]}\bar{f}_{i(t)}$. It follows from the definition of $S^{*}_{k}$ and $T(S^{*}_{k})$ that $\bar{f}_{i}\leq\bar{f}_{i(k)}$ for each $i\in[\eta]\setminus(S^{*}_{k}\cup T(S^{*}_{k}))$. Hence, it holds that
$\bar{f}_{i}\leq\frac{\sigma}{k}\text{ for all }i\in[\eta]\setminus(S^{*}_{k}\cup T(S^{*}_{k})).$ (24)
One may easily check that $(S^{*}\setminus S^{*}_{k})\cap T(S^{*}_{k})=\emptyset$. Thus, $(y^{*},\alpha^{*})$ is feasible to $LP(S^{*}_{k})$, and so
$\zeta^{*}\geq\sum_{i\in[\eta]}\bar{f}_{i}y^{\prime}_{i}(S^{*}_{k})+\sum_{j\in[\mu]}\bar{v}_{j}\alpha^{\prime}_{j}(S^{*}_{k})\geq\sum_{i\in[\eta]}\bar{f}_{i}y^{\prime\prime}_{i}(S^{*}_{k})-\delta+\sum_{j\in[\mu]}\bar{v}_{j}\alpha^{\prime}_{j}(S^{*}_{k})\geq\zeta-\delta,$
where $\delta=\sum_{i\in D}\bar{f}_{i}$ and $D=\{i\in[\eta]\mathrel{\mathop{\ordinarycolon}}0<y^{\prime}_{i}(S^{*}_{k})<y^{\prime\prime}_{i}(S^{*}_{k})\}$. Note that every $i\in D$ implies $y_{i}$ is a basic variable in $y^{\prime}(S^{*}_{k})$. Thus $|D|\leq\mu$. Moreover, $D\cap(S^{*}_{k}\cup T(S^{*}_{k}))=\emptyset$, and so we have $\bar{f}_{i}\leq\sigma/k$ for all $i\in D$ due to (24). Hence $\delta\leq\mu\sigma/k$, and $\zeta^{*}\geq\zeta-\mu\sigma/k\geq\zeta-\mu\zeta^{*}/k$ since $\zeta^{*}\geq\sigma$. Therefore, $\zeta\leq(1+\mu/k)\zeta^{*}\leq(1+\epsilon)\zeta^{*}.$
The algorithm’s runtime is dominated by solving the linear program on line 10 for each of the $\mathcal{O}(\eta^{k})$ subsets $S$. cohenSolvingLinearPrograms2021 present the current state-of-the-art method to solve a linear program in $\mathcal{O}\left(M(\eta)\log\eta\log\left(\eta/2^{-\mathcal{O}(L)}\right)\right)$ time where $M(\eta)\sim\eta^{2.38}$ is the cost of matrix multiplication and inversion, and $L=\mathcal{O}\left(\log(\eta+\mathinner{\!\left\lVert\bar{d}\right\rVert}_{\infty}+\mathinner{\!\left\lVert\bar{f}\right\rVert}_{\infty}+\mathinner{\!\left\lVert\bar{v}\right\rVert}_{\infty})\right)$. A consequence of using their method is that while optimal primal and dual solutions are obtained, an optimal basis is not necessarily identified. Therefore, we apply the algorithm due to megiddoFindingPrimalDualOptimal1991 to compute an optimal basis in strongly polynomial time from an optimal primal-dual solution pair. ∎
Theorem 3.2.
There exists a polynomial-time approximation scheme for p when $m$ is fixed.
Proof.
Proposition 2.4 shows a decomposition of p into $\mathcal{O}(n^{m})$ instances of the multidimensional knapsack cover problem with a single continuous variable per dimension. One can run Algorithm 1 on each of these instances and output the cheapest solution. By Corollary 2.5 and Theorem 3.1, this leads to a PTAS for p when $m$ is fixed. ∎
Theorem 3.2 gives, in a sense, the best polynomial-time approximation one can design to p since the multidimensional knapsack cover problem – even without continuous variables – does not admit any fully polynomial-time approximation scheme unless $\mathsf{P}=\mathsf{NP}$ (magazine1984note).
3.1 Extension to packing problems
A PTAS for the packing variant of p can be derived with minor modifications to Algorithm 1, as shown in Algorithm 2. The key changes are threefold. First, $\bar{d}_{j}(S)$ is redefined as $\bar{d}_{j}(S)=\bar{d}_{j}-\bar{\ell}_{j}-\sum_{i\in S}w_{ij}$ for each $j\in[\mu]$. Second, the feasibility check becomes verifying that $\bar{d}_{j}(S)\geq 0$ for each $j\in[\mu]$. Finally, after solving the linear program $LP(S)$, its solutions must be rounded down.
Algorithm 2 Algorithm $\mathcal{A}_{\epsilon}$ for mkp.
1:An instance $I=(\eta,\mu,\bar{f},\bar{v},\bar{c},\bar{w},\bar{d})$ of mkp.
2:A feasible solution $(y,\alpha)\in\{0,1\}^{\eta}\times\mathbb{R}^{\mu}$ to $I$.
3:
4:procedure $\mathcal{A}_{\epsilon}$(II)
5:  $k$ $\leftarrow$ $\min\{\eta,\>\lceil\mu/\epsilon\rceil\}$
6:  $\zeta$ $\leftarrow$ 0
7:  for $S\subseteq[\eta]$ with $|S|\leq k$ do
8:   Let $\bar{d}_{j}(S)\mathrel{\mathop{\ordinarycolon}}=\bar{d}_{j}-\bar{\ell}_{j}-\sum_{i\in S}w_{ij}$ for each $j\in[\mu]$
9:   if $\bar{d}_{j}(S)\geq 0$ for each $j\in[\mu]$ then $\triangleright$ Otherwise, $LP(S)$ is infeasible
10:     Compute an optimal basic solution $(y^{\prime}(S),\alpha^{\prime}(S))$ to $LP(S)$
11:     $y^{\prime\prime}_{i}(S)$ $\leftarrow$ $\left\lfloor y^{\prime}_{i}(S)\right\rfloor$ for every $i\in[\eta]$ $\triangleright$ Rounding down to an integer solution
12:     $\zeta(S)$ $\leftarrow$ $\sum_{i\in[\eta]}f_{i}y^{\prime\prime}_{i}(S)+\sum_{j\in[\mu]}\bar{v}_{j}\alpha^{\prime}_{j}(S)$
13:     if $\zeta<\zeta(S)$ then
14:       $\zeta$ $\leftarrow$ $\zeta(S)$,   $y$ $\leftarrow$ $y^{\prime\prime}(S)$,   $\alpha$ $\leftarrow$ $\alpha^{\prime}(S)$
15:  return $(y,\alpha)$
3.2 Extension to assignment problems
The approach previously described can also be extended to yield a PTAS for a related class of problems of the following form, which we denote by q:
$\max\;\;$ $\sum_{i\in[n]}\sum_{j\in[m]}v_{ij}x_{ij}+\sum_{i\in[n]}\sum_{j\in[m]}f_{ij}y_{ij}$
s.t. $\sum_{i\in[n]}x_{ij}\leq d_{j}$ $\forall j\in[m],$
$\sum_{j\in[m]}y_{ij}\leq 1$ $\forall i\in[n],$
$\ell_{ij}y_{ij}\leq x_{ij}\leq c_{ij}y_{ij}$ $\forall i\in[n],\;j\in[m],$
$y_{ij}\in\{0,1\}$ $\forall i\in[n],\;j\in[m],$
where $m$ and $n$ are positive integers, $v,\ell,c,f\in\mathbb{Z}_{\geq}^{nm}$ and $d\in\mathbb{Z}_{\geq}^{m}$ denote the nonnegative integer parameters. It is assumed, without loss of generality, that $d_{j}\geq c_{ij}\geq\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$. Note that q includes relevant problems as the Generalized Assignment Problem (gap). In particular, when $v_{ij}=0$ and $c_{ij}=\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$q is equivalent to gap, which is known to be $\mathsf{APX}$-hard (chekuri2005polynomial). On the positive side, for any $\epsilon>0$, there is an approximation for gap with ratio $(e/(e-1)+\epsilon)\approx(1.582+\epsilon)$ using an LP-rounding approach, which is currently the best known approximation for this problem (fleischerTightApproximationAlgorithms2006).
For a fixed $m$, the results from Section 2 can be adapted to decompose q into $\mathcal{O}(n^{m})$ instances of gap with a single continuous variable for each dimension as follows. Define $\bar{n}=n+m$ by introducing $m$ “dummy” items $\{n+1,n+2,\ldots,n+m\}$. For each $i\in[\bar{n}]\setminus[n]$ and $j\in[m]$, we set $\ell_{ij}=c_{ij}=v_{ij}=f_{ij}=0$. These newly created items ensure the existence of solutions where no item $i\in[n]$ is selected for a certain $j\in[m]$. The main difference compared to the approach in Section 2 lies in the definition of $\mathcal{G}$, here it is the set of all injective functions from $[m]$ to $\bar{n}]$. For each $g\in\mathcal{G}$, the formulation for $g$ is:
$\max\;\;$ $\sum_{i\in[\bar{n}]}\sum_{j\in[m]}f^{g}_{ij}y^{g}_{ij}+\sum_{j\in[m]}v^{g}_{j}\alpha^{g}_{j}$
s.t. $\sum_{i\in[\bar{n}]}w^{g}_{ij}y^{g}_{ij}+\alpha^{g}_{j}\leq d^{g}_{j}$ $\forall j\in[m],$
$\sum_{j\in[m]}y^{g}_{ij}\leq 1$ $\forall i\in[\bar{n}],$
$0\leq\alpha^{g}_{j}\leq c^{g}_{g(j),j}$ $\forall j\in[m],$
$y^{g}_{g(j),j}=1$ $\forall j\in[m],$
$y^{g}_{ij}\in\{0,1\}$ $\forall i\in[\bar{n}],j\in[m],$
where $w^{g}_{ij},f^{g}_{ij},v^{g}_{j},d^{g}_{j},c^{g}_{g(j),j}$ are defined in the same way as in Section 2.
Because $m$ is assumed to be fixed, we can transform this gap instance with one continuous variable per dimension into an instance of mkp with some additional cardinality constraints. We define a new set of $mn+m$ items as follows. For each original item $i\in[n]$ and each $j\in[m]$, we create a new item that has weight $w^{g}_{ij}$ and profit $f^{g}_{ij}$ for knapsack $j$, and a weight and profit equal to 0 for all other knapsacks $j^{\prime}\in[m]\setminus\{j\}$. We retain the $m$ dummy items from the previous augmentation, each with weight 0 and profit 0 for all knapsacks. We additionally add a cardinality constraint for each $i\in[n]$ to mkp, namely $\sum_{j\in[m]}y^{g}_{ij}\leq 1$ for all $i\in[n]$. Note that no cardinality constraints are added for the dummy items. We then apply the PTAS presented in Algorithm 2 to this new instance. The algorithm remains unchanged, except that in the feasibility check (Step 9 of Algorithm 2), it additionally ensures that at most one copy of each item $i\in[n]$ is selected. This adjustment does not affect the algorithm’s analysis. This leads us to the following analogous version of Theorem 3.2 for q:
Theorem 3.3.
There exists a polynomial-time approximation scheme for q when $m$ is fixed.
Furthermore, this approach also yields a PTAS for the covering version of q. Moreover, by replacing $\sum_{j\in[m]}y_{ij}\leq 1$ with $\sum_{j\in[m]}y_{ij}\leq t_{i}$ for all $i\in[n]$ for a parameter $t\in[n]^{n}$, we can design a PTAS for a family of related problems using the same core method. Observe that we obtain q when $t_{i}=1$ for every $i\in[n]$, and the packing version of p when $t_{i}=n$ for all $i\in[n]$.
4 FPTAS and approximate formulation for $m=1$
In this section, we design a fully polynomial-time algorithm for p when $m=1$, denoted as p1. This problem generalizes the knapsack cover problem as this occurs when $\ell_{i}=c_{i}$ for all $i\in[n]$. Hence p1 is clearly $\mathsf{NP}$-hard. Theorem 3.2 already proves that a PTAS exists for this problem. From the introduction we also know that vanhoeselFullyPolynomialApproximation2001 present an FPTAS for the single-item capacitated economic lot-sizing problem (which includes the knapsack cover problem) that is more general than p1 in its consideration of backlogging and holding costs. However, their model does not include lower bounds. This distinction in problem scope is accompanied by a methodological divergence as their approach does not rely on a decomposition into knapsack sets. The current state-of-the-art approximation algorithm for the knapsack problem is a randomized pseudo-polynomial algorithm by qizheng2024simple, which achieves a time complexity of $\mathcal{O}(n^{3/2}\times\min\{w_{\max},p_{\max}\})$. However, it is not obvious how the introduction of a continuous variable affects this result, a question that warrants its own investigation.
The formulation presented in Proposition 2.4 reveals a decomposition of p1 into a collection of $\mathcal{O}(n)$ instances of the knapsack cover problem with a single continuous variable, denoted by 1-mkc. Consider an arbitrary (but fixed) $g\in[n]$. As observed in Section 2, problem p1 for $g$ is equivalent to an instance $(\eta,\bar{v},\bar{f},\bar{w},\bar{d},\bar{c})$ of 1-mkc, which can be modeled as
$\min\;\;$ $\sum_{i\in[\eta]}\bar{f}_{i}y_{i}+\bar{v}\alpha$
s.t. $\sum_{i\in[\eta]}\bar{w}_{i}y_{i}\geq\bar{d}-\alpha,$
$\bar{c}\geq\alpha\geq 0,$
$y_{i}\in\{0,1\}$ $\forall i\in[\eta].$
Note that Corollary 2.5 can also be used in algorithms to solve p1. In what follows, we propose an FPTAS for 1-mkc which is outlined in Algorithm 3. The core of this algorithm is an exact approach that runs in pseudo-polynomial time as outlined in Procedure LABEL:alg:name:dp of Algorithm 3. This is a dynamic programming algorithm for 1-mkc without a continuous variable which is essentially the same as the folklore DP for the classic knapsack (packing) problem. For the sake of simplicity, this algorithm computes only the value of the solution, however one may easily modify it to store the solution itself. Precisely, procedure LABEL:alg:name:dp computes, for each $i\in[\eta]$ and $j\in[\sum_{i\in[\eta]}\bar{f}_{i}]$:
$M(i,j)=\max\left\{\sum_{s\in S}\bar{w}_{s}\mathrel{\mathop{\ordinarycolon}}S\subseteq\{1,\ldots,i\}\;\text{and}\;\sum_{s\in S}\bar{f}_{s}\leq j\right\}.$
Intuitively, $M$ stands for the maximum amount that can be covered at each cost level $j$ using a subset of the first $i$ elements in $\eta]$. After computing $M$, Algorithm 3 proceeds to compute the optimal value of the continuous variable.
Algorithm 3 Algorithm $\mathcal{A}^{1}_{\epsilon}$ for 1-mkc.
1:An instance $I=(\eta,\bar{v},\bar{f},\bar{w},\bar{d},\bar{c})$ of 1-mkc.
2:The value of a feasible solution $(y,\alpha)\in\{0,1\}^{\eta}\times\mathbb{R}$ of $I$.
3:
4:procedure $\mathcal{A}^{1}_{\epsilon}$(II)
5:  $I^{\prime}=(\eta,v^{\prime},f^{\prime},\bar{w},\bar{d},\bar{c})$ $\leftarrow$ LABEL:alg:name:scaling(II)
6:  $M$ $\leftarrow$ LABEL:alg:name:dp($I^{\prime}$)
7:  $q$ $\leftarrow$ $\infty$
8:  $j=\sum_{i\in[\eta]}f^{\prime}_{i}$
9:  while $M(\eta,j)\geq\bar{d}-\bar{c}$ and $j\geq 0$ do
10:   if $j+v^{\prime}(\max\{\bar{d}-M(\eta,j),0\})<q$ then
11:     $q=j+v^{\prime}(\max\{\bar{d}-M(\eta,j),0\})$
12:     j=j1j=j-1
13:  $(y,\alpha)$ $\leftarrow$ A feasible solution to $I^{\prime}$ of value $q$
14:  return $(y,\alpha)$
15:
16:procedure Scaling(II)
17:  $f_{\max}\leftarrow\max_{i\in[\eta]}\bar{f}_{i}$
18:  $\lambda\leftarrow\max\{\epsilon f_{\max}/(\eta\operatorname{poly}(\eta));\>1\}$
19:  for $i\in[\eta]$ do
20:   $f^{\prime}_{i}\leftarrow\left\lceil\bar{f}_{i}/{\lambda}\right\rceil$
21:  $v^{\prime}$ $\leftarrow$ $\bar{v}/\lambda$
22:  return $(\eta,v^{\prime},f^{\prime},\bar{w},\bar{d},\bar{c})$
23:
24:procedure DP(II)
25:  $M(i,0)=0$ for all $i\in[\eta]$
26:  for $i\in[\eta]$ do
27:   for $j$ from 0 to $\sum_{i\in\eta}\bar{f}_{i}$ do
28:     if $\bar{f}_{i}>j$ then
29:      $M(i,j)$ $\leftarrow$ M(i1,j)M(i-1,j)
30:     else
31:      $M(i,j)$ $\leftarrow$ $\max\{M(i-1,j);\>\bar{w}_{i}+M(i-1,j-\bar{f}_{i})\}$
32:  return $M$
Theorem 4.1.
Let $I=(\eta,\bar{v},\bar{f},\bar{w},\bar{d},\bar{c})$ be an instance of 1-mkc with $f_{\max}\geq f_{\min}\geq f_{\max}/\operatorname{poly}(\eta)$, where $f_{\max}=\max_{i\in[\eta]}\bar{f}_{i}$ and $f_{\min}=\min\left\{\min_{i\in[\eta]}\bar{f}_{i},\>\bar{v}\bar{d}\right\}$, and $\operatorname{poly}(\eta)\geq 1$ is a polynomial function on $\eta$. Algorithm 3 is a $(1+\epsilon)$-approximation that runs in $\mathcal{O}\left(\eta^{3}\operatorname{poly}(\eta)/\epsilon\right)$ time, for any $\epsilon>0$.
Proof.
We first analyze the algorithm supposing that no scaling is performed in line 5, that is, $I^{\prime}=I$. Let $q$ be the value produced by $\mathcal{A}^{1}_{\epsilon}$ on input $I$. It follows from the computation of $M$ by procedure LABEL:alg:name:dp and the conditions in line 9 that $q$ corresponds to the value of a feasible solution of $I$. Let $(y^{*},\alpha^{*})$ be an optimal solution for $I$ with $\alpha^{*}$ integer. The existence of this solution is guaranteed since $\bar{w},\bar{d}$ and $\bar{c}$ are all integer. Let $j^{*}=\sum_{i\in[\eta]}\bar{f}_{i}y^{*}_{i}$. Note that $M(\eta,j^{*})\geq\sum_{i\in[\eta]}\bar{w}_{i}y^{*}_{i}\geq\bar{d}-\alpha^{*}$. Hence, $\alpha^{*}\geq\bar{d}-M(n,j^{*})$. Therefore, we have $q\leq j^{*}+\bar{v}(\bar{d}-M(n,j^{*}))\leq\sum_{i\in[\eta]}\bar{f}_{i}y^{*}_{i}+\bar{v}\alpha^{*}=\operatorname{OPT}(I)$. Clearly, the running time of this exact algorithm depends on $\bar{f}$. This explains why we first scale down the input using procedure LABEL:alg:name:scaling in line 5 before executing the pseudopolynomial-time algorithm to solve the problem. As we shall see, this transformation guarantees that the algorithm runs in polynomial time on the size of $I$ if $f^{\max}/f^{\min}$ is bounded by a polynomial function on $\eta$. Let $I^{\prime}$ denote the transformed instance in line 5. Let $q$ be the output of Algorithm 3 on input $I$, and let $(y,\alpha)\in\{0,1\}^{\eta}\times\mathbb{R}$ be a feasible solution of $I$ of value $q$. By the construction, $(y,\alpha)$ is optimal for $I^{\prime}$. Note that if $\lambda=1$, then this solution is also optimal for $I$. Hereafter, assume that $\lambda=\epsilon f_{\max}/(\eta\operatorname{poly}(\eta))>1$. We define $Q=\{i\in[\eta]\mathrel{\mathop{\ordinarycolon}}y_{i}=1\}$ and $Q^{*}=\{i\in[\eta]\mathrel{\mathop{\ordinarycolon}}y^{*}_{i}=1\}$. First, note that $(y^{*},\alpha^{*})$ is also a feasible solution for $I^{\prime}$, and so $\sum_{i\in Q^{*}}f^{\prime}_{i}+v^{\prime}\alpha^{*}\geq\sum_{i\in Q}f^{\prime}_{i}+v^{\prime}\alpha=\operatorname{OPT}(I^{\prime})$. It is clear from lines 1821 that $\frac{\bar{f}_{i}}{\lambda}\leq f^{\prime}_{i}\leq\frac{\bar{f}_{i}}{\lambda}+1$ for every $i\in[\eta]$ and $\frac{\bar{v}}{\lambda}=v^{\prime}$. Hence, the following sequence of inequalities hold for the solution $(y,\alpha)$ produced by Algorithm 3:
$\sum_{i\in Q}\bar{f}_{i}+\bar{v}\alpha$ $\leq\lambda\sum_{i\in Q}f^{\prime}_{i}+\lambda v^{\prime}\alpha=\lambda\operatorname{OPT}(I^{\prime})$
$\leq\lambda\sum_{i\in Q^{*}}f^{\prime}_{i}+\lambda v^{\prime}\alpha^{*}\leq\lambda\sum_{i\in Q^{*}}\left(\frac{\bar{f}_{i}}{\lambda}+1\right)+\lambda\left(\frac{\bar{v}}{\lambda}\right)\alpha^{*}$
$=\operatorname{OPT}(I)+\lambda|Q^{*}|\leq\operatorname{OPT}(I)+\dfrac{\epsilon f_{\max}}{\operatorname{poly}(\eta)}$
$\leq(1+\epsilon)\operatorname{OPT}(I),$
where the last inequality holds because $\operatorname{OPT}(I)\geq f_{\min}\geq f_{\max}/\operatorname{poly}(\eta)$.
The running-time of Algorithm 3 on input $I^{\prime}$ takes $\mathcal{O}(n^{2}f^{\prime}_{\max})$ time , with $f^{\prime}_{\max}=\max_{i\in[\eta]}f^{\prime}_{i}$. It follows from the scaling in lines 18-20 that $f^{\prime}_{\max}\leq f_{\max}/\lambda+1\leq\eta\operatorname{poly}(\eta)/\epsilon+1$. Therefore, Algorithm 3 runs in $\mathcal{O}\left(\left(\eta\operatorname{poly}(\eta)/\epsilon+1\right)\eta^{2}\right)$ time. ∎
The hypothesis in Theorem 4.1 on the polynomial bound on the ratio $f_{\max}/f_{\min}$ is due to fact that the optimal value of an instance of 1-mkc can be arbitrarily smaller than $f_{\max}$ in general. As an example, consider an instance of 1-mkc with $\eta=2$, $\bar{d}=1$, $f_{1}=2$, $w_{1}=1$, $f_{2}=100$, $w_{2}=100$, and $\bar{v}=100$. Here, $f_{\max}=100$ while the optimal value equals $2$. A key direction for future research is to eliminate the assumption on the polynomial bound on the ratio $f_{\max}/f_{\min}$. We hypothesize that an approach similar to that of bienstockTighteningSimpleMixedinteger2012 could be adapted for this purpose, but formally establishing this remains an open challenge.
For problem p1, Theorem 4.1 yields the following result.
Theorem 4.2.
There exists a fully polynomial-time approximation scheme for p when $m=1$, assuming $f_{\max}\geq f_{\min}\geq f_{\max}/\operatorname{poly}(n)$, where $f_{\max}=\max_{i\in n}f_{i}+c_{i}v_{i}$ and $f_{\min}=\min_{i\in n}f_{i}+\ell_{i}v_{i}$, and $\operatorname{poly}(n)\geq 1$ is any polynomial function on $n$.
Proof.
Proposition 2.4 shows a decomposition of p1 into $\mathcal{O}(n)$ instances of the knapsack cover problem with a single continuous variable. One can run Algorithm 3 on each of these instances and output the cheapest solution. By Corollary 2.5 and Theorem 4.1, this leads to an FPTAS for p1 when for every resulting knapsack cover problem with a single continuous variable, $f_{\max}\geq f_{\min}\geq f_{\max}/\operatorname{poly}(n)$, where $f_{\max}=\max_{i\in[\eta]}\bar{f}_{i}$ and $f_{\min}=\min\left\{\min_{i\in[\eta]}\bar{f}_{i},\>\bar{v}\bar{d}\right\}$, and $\operatorname{poly}(\eta)\geq 1$ is any polynomial function on $\eta$. It is clear that this is guaranteed by the assumption on the input data of p. The total running time of this algorithm is $\mathcal{O}(n^{4}\operatorname{poly}(n)/\epsilon)$. ∎
By utilizing a similar algorithm as $\mathcal{A}^{1}_{\epsilon}$, it is possible to obtain an analogous approximation result for the packing version of p1. Furthermore, in the packing case, the assumption of a polynomially bounded ratio $f_{\max}/f_{\min}$ is unnecessary, as the optimal value is at most $nf_{\max}$.
An important consequence of Theorem 3.2 and the equivalence of optimization and separation is that an $\epsilon$-approximate formulation must exist for p1. Since Proposition 2.4 decomposes p1 into $\mathcal{O}(n)$ instances of 1-mkc, we can directly leverage the results of bienstockTighteningSimpleMixedinteger2012, who obtain a provably tight, polynomially large formulation for the standard knapsack cover problem. We extend their methodology to incorporate a continuous variable, defining a linear programming formulation whose optimal value is at most $(1+\epsilon)$ times the optimal value of 1-mkc.
Leveraging this building block, we immediately obtain the following result for p1.
Theorem 4.3.
Let $\epsilon\in(0,1)$, and let $I$ be an instance of p1. There is a linear programming formulation $A^{I}z\leq b^{I}$ with $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}n^{3}\right)$ variables and constraints, and objective function $h^{I}$ such that
$\operatorname{OPT}(I)\leq(1+\epsilon)\min\{h^{I}(z)\mathrel{\mathop{\ordinarycolon}}A^{I}z\leq b^{I}\}.$
Proof sketch.
Proposition 2.4 shows a decomposition of p1 into $\mathcal{O}(n)$ instances of the knapsack cover problem with a single continuous variable. For each of these problems, one may create an $\epsilon$-approximate formulation with $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}n^{2}\right)$ variables and constraints, and combine them as in the formulation presented in Section 2. This leads to an $\epsilon$-approximate formulation for p1 with $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}n^{3}\right)$ variables and constraints. ∎
The proof of the $\epsilon$-approximate formulation for the knapsack cover problem with a single continuous variable can be found in the appendix as the approach and structure adheres closely to what is shown by bienstockTighteningSimpleMixedinteger2012.
The existence of an $\epsilon$-approximate formulation for the packing version of p1 is guaranteed in principle by the equivalence of optimization and separation. However, such a formulation is not immediately available. A key complication arises from the presence of a continuous variable after the decomposition of p1. While bienstockApproximateFormulations012008 successfully derived an approximate formulation for the standard 0-1 knapsack set, it is unclear how to adapt the techniques to handle an additional continuous variable. Resolving this gap, and thus developing an explicit $\epsilon$-approximate formulation for the packing variant of p1, is a significant open challenge.
5 Perfect compact formulation for $m=1$ and uniform bounds
In this section, we present a compact and tight formulation for p when $m=1$ and the bounds are uniform, that is, $\ell_{i}=\ell$ and $c_{i}=c$ for all $i\in[n]$. This special case is denoted as pu. As mentioned in the literature review, pu can be solved in polynomial time (hellionPolynomialTimeAlgorithm2012a). However, no tight formulation of the problem is known. Hence the proposed perfect compact formulation contributes in this direction.
constantinoLowerBoundsLotSizing1998a shows that solving pu, where $\ell_{i}$ can vary with $i\in[n]$, is in general $\mathsf{NP}$-hard. This paper also analyzes a setting identical to pu and obtains two families of valid inequalities. Using both families, it is possible to describe the convex hull when $c\geq d$. agraPolyhedralDescriptionInteger2006a study the case where the $y$ variables are general integers and unbounded.
dashContinuousKnapsackSet2016a study the convex hull of the continuous knapsack set. Among other results, they prove that the convex hull with two integer and one bounded continuous variable is described using inequalities for a knapsack set with two integer variables and the ones for a continuous knapsack set with two integer and one continuous variable that is unbounded. We use a special case of this result below.
We first present a result on the the convex hull of a simpler set. For two integers $a$ and $b$, we use $[a,b]$ to denote the set of integers $\{a,\ldots,b\}$ if $a\leq b$; otherwise $a,b]=\emptyset$.
Lemma 5.1.
Let
$Y=\{(\alpha,\psi)\in\mathbb{R}_{\geq}\times\{0,1\}^{\nu}\mathrel{\mathop{\ordinarycolon}}\alpha+\sum_{i\in[\nu]}\psi_{i}\geq\delta,\alpha\leq\sigma\},$
where $\delta$ and $\nu$ are positive, $\nu$ is an integer such that $\nu\geq\lceil\delta-\sigma\rceil$ and $\sigma\leq 1$. The description of the convex hull of $Y$ is given by the trivial inequalities,
$\sum_{i\in[\nu]}\psi_{i}\geq\lceil\delta-\sigma\rceil,\text{ and }$ (25)
$\alpha\geq(\delta-\lfloor\delta\rfloor)\left(\lceil\delta\rceil-\sum_{i\in[\nu]}\psi_{i}\right).$ (26)
Proof.
By Theorem 4.4. in dashContinuousKnapsackSet2016a, we know that $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}_{\geq}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\alpha+\gamma\geq\delta,\alpha\leq\sigma\}$ is the intersection of three sets, namely, $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}_{\geq}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\alpha+\gamma\geq\delta\}$, $\{(\alpha,\gamma)\in\mathbb{R}\times\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\alpha\leq\sigma\}$ and $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\gamma\geq\delta-\sigma\}$. In the nontrivial case, the mixed integer rounding inequality $\alpha\geq(\delta-\lfloor\delta\rfloor)(\lceil\delta\rceil-\gamma)$ is needed to describe the first set, and the integer rounding inequality $\gamma\geq\lceil\delta-\sigma\rceil$ is needed to describe the third set (see, e.g., wolsey2020integer). Adding an upper bound of $\nu$ on the integer variable $\gamma$ does not create fractional extreme points.
Now we know that in every extreme point of the set defined by the trivial inequalities and the valid inequalities (26) and (25), $\sum_{i=1}^{\nu}\psi_{i}$ is integer. If $\psi$ is fractional, then it has at least two entries that are fractional and such a point cannot be an extreme point. ∎
A perfect formulation for pu is obtained by integrating the decomposition into knapsack cover problems with one continuous variable presented in Section 2, and the union-of-polyhedra approach to the convex hull of sets described in Lemma 5.1.
Theorem 5.2.
pu admits the following perfect formulation with $\mathcal{O}(n^{3})$ variables and $\mathcal{O}(n^{3})$ constraints:
$\min\;\;$ $\sum_{i\in[n]}v_{i}x_{i}+\sum_{i\in[n]}f_{i}y_{i}$
s.t. $x_{i}=\sum_{g=i+1}^{n}\sum_{b\in B^{g}}\ell y_{i}^{gb}+\sum_{b\in B^{i}}x_{i}^{ib}+\sum_{g=1}^{i-1}\sum_{b\in B^{g}}cy_{i}^{gb}$ $\forall i\in[n],$
$y_{i}=\sum_{g\in\mathcal{G}}\sum_{b\in B^{g}}y_{i}^{gb}$ $\forall i\in[n],$
$\sum_{i=g+1}^{n}y^{gb}_{i}\geq\left\lceil\frac{d-b\ell-c}{c}\right\rceil z^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
$x_{g}^{gb}+\sum_{i=g+1}^{n}\left(d-(b+1)\ell-\left\lfloor\frac{(d-(b+1)\ell}{c}\right\rfloor c\right)y^{gb}_{i}$
$\hskip 28.45274pt\geq\left(\ell+\left\lceil\frac{d-(b+1)\ell}{c}\right\rceil\times\left(d-(b+1)\ell-\left\lfloor\frac{(d-(b+1)\ell}{c}\right\rfloor c\right)\right)z^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
$x_{g}^{gb}+\sum_{i=g+1}^{n}cy^{gb}_{i}\geq(d-b\ell)z^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
$\sum_{g\in\mathcal{G}}\sum_{b\in B^{g}}z^{gb}=1,$
$\sum_{i=1}^{g-1}y_{i}^{gb}=bz^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
$y_{g}^{gb}=z^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
$x_{g}^{gb}\geq\ell z^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
$x_{g}^{gb}\leq cz^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
$y_{i}^{gb}\leq z^{gb}$ $\forall i\in[n],g\in\mathcal{G},b\in B^{g},$
$y_{i}^{gb}\geq 0$ $\forall i\in[n],g\in\mathcal{G},b\in B^{g},$
$z^{gb}\geq 0$ $\forall g\in\mathcal{G},b\in B^{g},$
where $[n]$ is ordered so that $v_{1}\geq v_{2}\geq\ldots\geq v_{n}$, $\mathcal{G}=\left[1,\min\left\{\left\lfloor\frac{(n+1)c-d-\ell}{c-\ell}\right\rfloor,n\right\}\right]$, $B^{g}=\emptyset$ for all $g\in[n]\setminus\mathcal{G}$ and $B^{g}=\left[\max\left\{0,\left\lceil\frac{(d-(n+1-g)c)}{\ell}\right\rceil\right\},g-1\right]$ for $g\in\mathcal{G}$.
Proof.
First, $\min\left\{\left\lfloor\frac{(n+1)c-d-\ell}{c-\ell}\right\rfloor,n\right\}$ is the largest index of a fractional variable that leads to a feasible solution, $\mathcal{G}=\left[1,\min\left\{\left\lfloor\frac{(n+1)c-d-\ell}{c-\ell}\right\rfloor,n\right\}\right]$ is the set of possible fractional variables, $\max\left\{0,\left\lceil\frac{(d-(n+1-g)c)}{\ell}\right\rceil\right\}$ is the minimum value of $\sum_{i=1}^{g-1}y_{i}$ that leads to a feasible solution and $B^{g}=\left[\max\left\{0,\left\lceil\frac{(d-(n+1-g)c)}{\ell}\right\rceil\right\},g-1\right]$ is the set of possible values for $\sum_{i=1}^{g-1}y_{i}$ at solutions of $X^{g}$ for $g\in\mathcal{G}$.
Note that
$\operatorname{conv}(X)=\operatorname{conv}\bigcup_{g\in\mathcal{G}}\bigcup_{b\in B^{g}}\operatorname{conv}\left(X^{g}\cap\left\{(x,y)\mathrel{\mathop{\ordinarycolon}}\sum_{i=1}^{g-1}y_{i}=b\right\}\right),$
where $X^{g}$ is the set of all points $(x,y)$ that satisfy
$x_{i}=\ell y_{i}\quad i\in[g-1],$
$x_{i}=cy_{i}\quad i\in[g+1,n],$
$\sum_{i=1}^{g-1}\ell y_{i}+x_{g}+\sum_{i=g+1}^{n}cy_{i}\geq d,$
$\ell\leq x_{g}\leq c,$
$y_{g}=1,$
$y_{i}\in\{0,1\}\quad\forall i\in[n].$
For $g\in\mathcal{G}$ and $b\in B^{g}$, we have
$\operatorname{conv}\left(X^{g}\cap\left\{(x,y)\in\mathbb{R}^{n}_{\geq}\times\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}\sum_{i=1}^{g-1}y_{i}=b\right\}\right)$ $=\operatorname{conv}(X^{g1}(b))\cap\operatorname{conv}(X^{g2}(b))$
$\phantom{=}\cap\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}x_{i}=\ell y_{i}\>\forall i\in[g-1],$
[…]
where
$X^{g1}(b)=\left\{(x,y)\in\mathbb{R}^{n}\times\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}\sum_{i=1}^{g-1}y_{i}=b,y_{g}=1\right\},$
and
$X^{g2}(b)=\left\{(x,y)\in\mathbb{R}^{n}_{\geq}\times\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}x_{g}+\sum_{i=g+1}^{n}cy_{i}\geq d-\ell b,\ell\leq x_{g}\leq c\right\}.$
The description of $\operatorname{conv}(X^{g1}(b))$ is given by
$\sum_{i=1}^{g-1}y_{i}=b,$
$y_{g}=1,$
$0\leq y_{i}\leq 1\quad\forall i\in[g-1],$
and the description of $\operatorname{conv}(X^{g2}(b))$ is given by
$x_{g}+\sum_{i=g+1}^{n}cy_{i}\geq d-b\ell,$
$\ell\leq x_{g}\leq c,$
$0\leq y_{i}\leq 1\quad\forall i\in[g+1,n],$
$\sum_{i=g+1}^{n}y_{i}\geq\left\lceil\frac{d-b\ell-c}{c}\right\rceil,$
$x_{g}\geq\ell+\left(\frac{d-(b+1)\ell}{c}-\left\lfloor\frac{d-(b+1)\ell}{c}\right\rfloor\right)\left(\left\lceil\frac{d-(b+1)\ell}{c}\right\rceil-\sum_{i=g+1}^{n}y_{i}\right),$
after setting $\nu=n-g$, $\delta=\frac{d-(b+1)\ell}{c}$, $\sigma=\frac{c-\ell}{c}$ and $\alpha=\frac{x_{g}-\ell}{c}$ in Lemma 5.1. Using the result on the convex hull of the union of polyhedra (conforti2008compact; balas1998disjunctive), we obtain the tight formulation. ∎
A perfect compact formulation can also be derived for the packing version of pu. This is achieved by using sets $\mathcal{G}=[n]$ and $B^{g}=\left[0,\min\{\left\lfloor d-\ell/\ell\right\rfloor,g\}\right]$ for $g\in\mathcal{G}$. The primary distinction from the covering case lies in the description of the convex hull $\operatorname{conv}(X^{g2}(b))$, which can be characterized using an approach analogous to Lemma 5.1.
6 Conclusion and further research
This paper presents approximation schemes for various mixed-integer problems with a fixed number of constraints. By analyzing the structure of the extremes points of the associated polytope of such problems, we decompose it into multidimensional knapsack cover instances with a single continuous variable per dimension. By leveraging approximation algorithms from the knapsack literature, we obtain a PTAS for p. Utilizing a similar methodology, an FPTAS and an $\epsilon$-approximate formulation is presented for the one-dimensional case. For the one-dimensional case with uniform bounds, we propose a perfect compact formulation. To obtain these results, some classical knapsack results are extended to incorporate continuous variables. These results are especially valuable as they provide a framework to design algorithms and tight formulations for similarly structured problems such as fixed-charge, facility location, and supplier selection problems. We also derive analogous results for the packing and assignment variants of p.
Several promising directions for future research emerge from this work. A natural first direction is to extend the results to more general versions of p with bounded integer variables and piecewise-linear objective functions. Also, obtaining an $\epsilon$-approximate formulation for p presents an interesting avenue for further exploration.
Appendix A Appendix
Define 1-mkc:
$\operatorname{OPT}^{Z}=\;\;$ $\min\bar{v}\alpha+\sum_{i\in[\eta]}\bar{f}_{i}y_{i}$
s.t. $\sum_{i\in[\eta]}\bar{w}_{i}y_{i}\geq\bar{d}-\alpha$
$\bar{c}\geq\alpha\geq 0$
$y_{i}\in\{0,1\}$ $\forall i\in[\eta],$
where $\bar{f}$ and $\bar{v}$ are both positive, $0<\bar{w}_{i}\leq\bar{d}$ for all $i\in[\eta]$, and $0<\bar{c}\leq\bar{d}$. Let $OPT^{*}$ be the value of the LP relaxation of 1-mkc. Assume, without loss of generality, that $\bar{f}_{1}\geq\bar{f}_{2}\geq\ldots\geq\bar{f}_{\eta}$. In this appendix, we prove the following theorem using an approach and structure that adheres closely to what is shown by bienstockTighteningSimpleMixedinteger2012.
Theorem A.1.
Let $\epsilon\in(0,1)$. There exists a linear programming relaxation of 1-mkc with $\mathcal{O}\left((\frac{1}{\epsilon})^{\mathcal{O}(\frac{1}{\epsilon^{2}})}\eta^{2}\right)$ variables and constraints whose optimal value $\operatorname{OPT}(\epsilon)$ satisfies $\operatorname{OPT}^{Z}<(1+\epsilon)\operatorname{OPT}(\epsilon)$.
Before proving this result, we first show the following lemma.
Lemma A.2.
Let $H\geq 1$ be an integer. Suppose $S^{\prime}\subseteq[\eta]$, and let $0\leq\bar{y_{i}}\leq 1$ $(i\in S^{\prime})$ be given values. Let $f^{\max}=\max_{i\in S^{\prime}}\{\bar{f}_{i}\}$, $f^{\min}=\min_{i\in S^{\prime}}\{\bar{f}_{i}\}$.
(a) Suppose first that
$\sum_{i\in S^{\prime}}\bar{y_{i}}=H$
Then there exist 0/1 values $\hat{y_{i}}$ ($i\in S^{\prime}$) such that
$\sum_{i\in S^{\prime}}\bar{w}_{i}\hat{y_{i}}\geq\sum_{i\in S^{\prime}}\bar{w}_{i}\bar{y_{i}}$ (27)
and
$\sum_{i\in S^{\prime}}\bar{f}_{i}\hat{y_{i}}\leq(1-\frac{1}{H}+\frac{f^{\max}}{Hf^{\min}})\sum_{i\in S^{\prime}}\bar{f}_{i}\bar{y_{i}}$
(b) Suppose next that
$\sum_{i\in S^{\prime}}\bar{y_{i}}\geq H$
Then there exist 0/1 values $\hat{y_{i}}$ ($i\in S^{\prime}$) satisfying (27) and
$\sum_{i\in S^{\prime}}\bar{f}_{i}\hat{y_{i}}\leq(1+\frac{f^{\max}}{Hf^{\min}})\sum_{i\in S^{\prime}}\bar{f}_{i}\bar{y_{i}}$
Proof.
(a) Let $\dot{y}$ be an extreme point solution to the linear program
$\min\;\;$ $\sum_{i\in S^{\prime}}\bar{f}_{i}y_{i}$
s.t. $\sum_{i\in S^{\prime}}\bar{w}_{i}y_{i}\geq\sum_{i\in S^{\prime}}\bar{w}_{i}\bar{y_{i}}$ (28)
$\sum_{i\in S^{\prime}}y_{i}=H$ (29)
$0\leq y_{i}\leq 1$ $\forall i\in S^{\prime}$
Because this linear program only has two constraints (excluding variable bounds), any basic feasible solution (extreme point) has at most two basic variables. Consequently, at most two of the values $\dot{y_{i}}$, $i\in S^{\prime}$, are fractional, but since $H$ is integral, either zero or exactly two $\dot{y_{i}}$ are fractional. Suppose there exist $j,k\in S^{\prime}$ with $j\neq k$ such that
$0<\dot{y_{j}}<1,\quad 0<\dot{y_{k}}<1,\quad\dot{y_{j}}+\dot{y_{k}}=1.$
Assume, without loss of generality, that $\bar{w}_{j}\geq\bar{w}_{k}$. Then we can set $\hat{y_{j}}=1$, $\hat{y_{k}}=0$, and $\hat{y_{i}}=\dot{y_{i}}$ for all $i\in S^{\prime}\setminus\{j,k\}$, thereby obtaining a 0/1 vector $\hat{y}$ which satisfies (28) while increasing the cost of $\dot{y}$ by at most
$\bar{f}_{j}-\bar{f}_{j}\dot{y_{j}}-\bar{f}_{k}\dot{y_{k}}=(\bar{f}_{j}-\bar{f}_{k})(1-\dot{y_{j}})\leq f^{\max}-f^{\min}.$
Hence we obtain
$\frac{\sum_{i\in S^{\prime}}\bar{f}_{i}\hat{y_{i}}-\sum_{i\in S^{\prime}}\bar{f}_{i}\dot{y}_{i}}{\sum_{i\in S^{\prime}}\bar{f}_{i}\dot{y}_{i}}\leq\frac{f^{\max}-f^{\min}}{\sum_{i\in S^{\prime}}\bar{f}_{i}\dot{y}_{i}}\leq\frac{f^{\max}-f^{\min}}{Hf^{\min}}.$
(b) Proceeding in a similar way to (a) (using, instead of (29), $\sum_{i\in S^{\prime}}y_{i}\geq H$), it can be assumed that either zero, one or two of the $\dot{y_{i}}$ ($i\in S^{\prime}$) are fractional. If two are fractional the result is implied by (a). If there is only one fractional $\dot{y_{i}}$ then rounding up $\dot{y}$ provides a 0/1 vector $\hat{y}$ that is feasible while increasing the cost by at most $f^{\max}$. Hence it holds that
$\frac{\sum_{i\in S^{\prime}}\bar{f}_{i}\hat{y_{i}}-\sum_{i\in S^{\prime}}\bar{f}_{i}\dot{y}_{i}}{\sum_{i\in S^{\prime}}\bar{f}_{i}\dot{y}_{i}}\leq\frac{f^{\max}}{Hf^{\min}}.$
Let $\epsilon\in(0,1)$. Define $K$ as the smallest integer such that $(1+\epsilon)^{-K}\leq\epsilon$. Note that $K\leq 1+\log(1/\epsilon)/\log(1+\epsilon)$. Let us define $J=\lceil 1+\frac{1}{\epsilon}\rceil$.
Definition A.3.
A signature $\sigma$ is a vector in $\mathbb{Z}^{K}$ such that $0\leq\sigma_{i}\leq J$ for $i\in[K]$.
For every $h\in[\eta]$ and $k\in[K]$, define
$S^{h,k}=\left\{i\in[\eta]\mathrel{\mathop{\ordinarycolon}}\bar{f}_{h}(1+\epsilon)^{-(k-1)}\geq\bar{f}_{i}>\bar{f}_{h}(1+\epsilon)^{-k}\text{ and }i>h\right\}.$
For each $h\in[\eta]$ and each signature $\sigma$, define
$\begin{split}P^{h,\sigma}=\Bigg\{(y,\alpha)\in[0,1]^{\eta}\times\mathbb{R}\mathrel{\mathop{\ordinarycolon}}&\sum_{i\in[\eta]}\bar{w}_{i}y_{i}\geq\bar{d}-\alpha,\quad\bar{c}\geq\alpha\geq 0,\\ &y_{1}=y_{2}=\ldots=y_{h-1}=0,\quad y_{h}=1,\\ &\sum_{i\in S^{h,k}}y_{i}=\sigma_{k}\quad\forall k\mathrel{\mathop{\ordinarycolon}}\sigma_{k}<J,\\ &\sum_{i\in S^{h,k}}y_{i}\geq J\quad\forall k\mathrel{\mathop{\ordinarycolon}}\sigma_{k}=J\Bigg\}\end{split}$
Note that the sets $S^{h,k}$ partition the variables $y_{i}$ whose cost $\bar{f}_{i}$ lies between $\bar{f}_{h}$ and $\epsilon\bar{f}_{h}$. Furthermore, all variables in each set $S^{h,k}$ have nearly the same cost. For a given $P^{h,\sigma}$, the signature $\sigma$ counts the number of $y_{i}$ that take value $1$ in each $S^{h,k}$ with $k\in[K]$. Thus every feasible solution for 1-mkc belongs to some $P^{h,\sigma}$. To see why, consider a feasible solution of 1-mkc and let $h$ be the smallest index $i\in[\eta]$ for which $y_{i}=1$. For any set $S^{h,k}$, let $\ell$ be the number of indices $i\in[\eta]$ within this set for which $y_{i}=1$. The value of $\sigma_{k}$ is then defined as $\min\{\ell,J\}$.
Lemma A.4.
For each $h\in[\eta]$ and signature $\sigma$ with $P^{h,\sigma}\neq\emptyset$, there is $(\hat{y}^{h,\sigma},\hat{\alpha}^{h,\sigma})\in\mathbb{B}^{\eta}\times\mathbb{R}_{\geq}$ which is feasible for 1-mkc such that $\bar{f}^{T}\hat{y}^{h,\sigma}+\bar{v}\hat{\alpha}^{h,\sigma}\leq(1+\epsilon)\min\{\bar{f}^{T}y+\bar{v}\alpha\mathrel{\mathop{\ordinarycolon}}(y,\alpha)\in P^{h,\sigma}\}$. As a result, $\operatorname{OPT}^{Z}\leq(1+\epsilon)\min\{\bar{f}^{T}y+\bar{v}\alpha\mathrel{\mathop{\ordinarycolon}}(y,\alpha)\in P^{h,\sigma}\}$.
Proof.
Let $(\bar{y},\bar{\alpha})\in P^{h,\sigma}$. Define $f^{\max}=\max_{i\in S^{h,k}}\{\bar{f}_{i}\}$, $f^{\min}=\min_{i\in S^{h,k}}\{\bar{f}_{i}\}$. Let $\hat{y}^{h,\sigma}$ be defined as follows. First, for each $k\in[K]$ such that $\sigma_{k}>0$, we obtain the values $\hat{y}_{i}^{h,\sigma}$ for each $i\in S^{h,k}$ by applying Lemma A.2 with $S^{\prime}=S^{h,k}$; note that when $\sigma_{k}<J$, then we have
$1-\frac{1}{\sigma_{k}}+\frac{f^{\max}}{\sigma_{k}f^{\min}}\leq\left(1-\frac{1}{\sigma_{k}}\right)\frac{f^{\max}}{f^{\min}}+\frac{f^{\max}}{\sigma_{k}f^{\min}}=\frac{f^{\max}}{f^{\min}}\leq 1+\epsilon$
by construction of the sets $S^{h,k}$. If $\sigma_{k}=J$, we have
$1+\frac{f^{\max}}{Jf^{\min}}\leq 1+\epsilon$
by our choice for $J$. If on the other hand $\sigma_{k}=0$ we set $\hat{y}_{i}^{h,\sigma}=0$ for every $i\in S^{h,k}$. Finally, define $T^{h}=\{i\in[\eta]\mathrel{\mathop{\ordinarycolon}}\bar{f}_{i}\leq(1+\epsilon)^{-K}\bar{f}_{h}\}$. Thus the $S^{h,k}$ for all $l\in[K]$, together with $T^{h}$, form a partition $\eta-h]$. The problem
$\min\;\;$ $\sum_{i\in T^{h}}\bar{f}_{i}y_{i}$
s.t. $\sum_{i\in T^{h}}\bar{w}_{i}y_{i}\geq\sum_{i\in T^{h}}\bar{w}_{i}\bar{y_{i}}$
$0\leq y_{i}\leq 1$ $\forall i\in T^{h}$
is a linear relaxation of a knapsack cover problem, and hence it has an optimal solution $y^{*}$ with at most one fractional variable. We set $\hat{y}_{i}^{h,\sigma}=\lceil y^{*}_{i}\rceil$ for each $i\in T^{h}$; thereby increasing cost (from $y^{*}$) by less than $(1+\epsilon)^{-K}\bar{f}_{h}\leq\epsilon\bar{f}_{h}$ by definition of $K$ and $\hat{\alpha}^{h,\sigma}=\bar{\alpha}$. In summary,
$\bar{f}^{T}\hat{y}^{h,\sigma}+\bar{v}\hat{\alpha}^{h,\sigma}-\bar{f}^{T}\bar{y}-\bar{v}\bar{\alpha}$ $=\sum_{k\mathrel{\mathop{\ordinarycolon}}\sigma_{k}>0}\left(\sum_{i\in S^{h,k}}\bar{f}_{i}\hat{y_{i}}^{h,\sigma}-\sum_{i\in S^{h,k}}\bar{f}_{i}\bar{y_{i}}\right)+\sum_{i\in T^{h}}\bar{f}_{i}(\hat{y}_{i}^{h,\sigma}-\bar{y}_{i})$
$\leq\epsilon\sum_{k\mathrel{\mathop{\ordinarycolon}}\sigma_{k}>0}\sum_{i\in S^{h,k}}\bar{f}_{i}\bar{y_{i}}+\epsilon\bar{f}_{h}$ (30)
$\leq\epsilon\bar{f}^{T}\bar{y}$ (31)
Here, the equality follows from $\hat{\alpha}^{h,\sigma}=\bar{\alpha}$, (30) follows from Lemma A.2 and by definition of sets $S^{h,k}$ and (31) follows from the fact that $\bar{y}_{h}=1$, by definition of $P^{h,\sigma}$. ∎
Consider the polyhedron
$Q=\operatorname{conv}\left(\bigcup_{h\in[\eta],\sigma\in([J]\cup\{0\})^{K}}P^{h,\sigma}\right)$
Note that there are at most $(J+1)^{K}\eta=\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}\eta\right)$ polyhedra $P^{h,\sigma}$, and that each $P^{h,\sigma}$ is described by a system with $\mathcal{O}(K+\eta)$ constraints in $\eta$ variables. Thus, $Q$ is the projection to $\mathbb{R}^{\eta+1}$ of the feasible set for a system with at most $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}\eta^{2}\right)$ constraints and variables. Furthermore, any 0/1 vector $y$ and $\alpha$ that is feasible for 1-mkc  satisfies $y\in P^{h,\sigma}$ for some $h$ and $\sigma$. In other words, $Q$ constitutes a valid relaxation to 1-mkc.
Lemma A.5.
$\operatorname{OPT}^{Z}\leq(1+\epsilon)\min\{\bar{f}^{T}y+\bar{v}\alpha\mathrel{\mathop{\ordinarycolon}}(y,\alpha)\in Q\}$.
Proof.
[…]